OpenStudy (anonymous):

f ''(θ) = sin θ + cos θ, f(0) = 2, f '(0) = 2 I will give medal to anyone who help! Thanks. For F(t) I got -sin(t) - cos(t) + t + 3 Did I do it right? Thanks! :))

OpenStudy (anonymous):

\[\begin{align*}f''(\theta)&=\sin\theta+\cos\theta\\ \int f''(\theta)~d\theta&=\int(\sin\theta+\cos\theta)~d\theta\\ f'(\theta)&=-\cos\theta+\sin\theta+C_1&(1)\\ \int f'(\theta)~d\theta&=\int(\sin\theta-\cos\theta+C_1)~d\theta\\ f(\theta)&=-\cos\theta-\sin\theta+C_1\theta+C_2&(2) \end{align*}\] Plug the initial value \(f'(0)=2\) into equation (1), so \[2=-\cos0+\sin0+C_1~~\Rightarrow~~C_1=3\] Plug the other initial value into eq (2): \[2=-\cos0-\sin0+3(0)+C_2~~\Rightarrow~~C_2=3\] Finalsolution should be \[f(\theta)=-\cos\theta-\sin\theta+3\theta+3\]